# High stakes exams inflate a gender gap and contribute to systematic grading errors in introductory physics

### v2

@@ I.1.1 Gaps - student level characteristics
Many different explanations for the existence of these types of gender gaps have been proposed and investigated. Some of the leading theories for the gender gaps in physics are that women on average i) have poorer physics and/or math preparation [9, 10, 6], ii) have less overall interest in physics [11], and/or iii) suffer more from test anxiety, which is amplified on high-stakes exams [12, 13, 14, 10, 15, 16, 7]. It’s also been shown that women don’t feel like they belong in physics as much as their male peers [17, 4], that they have less of a science or physics identity [18], and/or have lower self-efficacy in physics [10, 15, 19, 20].
@@ I.1.2 Gaps - system level issues
Leslie et al. [22] take a more systemic view in noting that gender gaps in representation are largest in a few specific academic fields (including physics) where the practitioners in the field are more likely to think that to succeed in their field “hard work alone just won’t cut it; you need to have an innate gift or talent.” Muradoglu et al. [23] use this idea, that success in a field requires innate brilliance, together with the general societal view that women are not as intellectually talented as men [24], to argue that a natural result is that, on average, women would have lower average self-efficacy in such fields. Since, self-efficacy inversely correlates with anxiety, it also is reasonable that women, on average, would be expected to be more anxious than men during the high stakes exams in these particular academic fields. As noted earlier, in this paper we will show data consistent with the idea that there is no gender gap in physics understanding/ability but that the commonly measured exam gender gap arises because of the stakes associated with that exam. In this sense the exam gender gap may simply be an incorrect measurement of the underlying knowledge/ability. It would be ironic if such illusory exam gender grade gaps were part of the evidence society used to conclude that women were not as talented as men at physics and that that conclusion led to women being more anxious on, and under-performing on, exactly these high stakes exams that led to society’s conclusions.
@@ I.2 Retake Exams in Physics
One possible way to lower the stakes of an exam, while still using exams as an assessment, is to offer a retake option for each exam given during the term. For each in-term exam, a retake covers the same topics as the first-attempt exam, is offered after the students have seen their score on their first-attempt, and is given a score that can supplant the first-attempt score. Thus far, there is very little research on retake exams in large introductory STEM classes. Davis et al. [26] find that students in a large introductory human physiology course who took advantage of retake exams on average increase their grade by 6%. Kortemeyer et al. [27] share the benefits of retakes as learning tools in introductory physics. However, none of the examples we found examined the relationship between the retakes and the gender gap in physics courses aside from our own previous work on this same dataset.
@@ III Database Details
All of the exam data we discuss here are from classes in the introductory physics series taken mostly by bioscience students and offered between 1997 and 2015 at an R1 public university. We have access to individual exam scores for three of the four classes that offered students retake exams. To compare the exam scores from the retake classes to the exam scores from the regular classes which did not offer retakes we add, to the database, exam data of the 254 non-retake classes from the same series for which we have scores from exams given to students during the term. We also added each student’s gender and race/ethnicity as well as each student’s university GPA upon entering the class. These data were obtained from the university administration and included before anonymizing the entire dataset several years ago. Since the numerical exam scores are given in a wide variety of ways by the many instructors of these classes, we have z-scored each exam in each course to give each exam grade distribution the same average, 0, and the same standard deviation, 1.
@@ IV.1 In-term exams in retake classes
Our expectation was that students who had high scores on their first attempt would not risk lowering that score and so would not take a retake exam. We found that on average a first attempt of an exam was followed by 34% of the students taking the retake a week later. The retake classes were anomalous in having zero gender gap in course grades (i.e. female students had the same average course grades as male students unlike in the regular courses) suggesting the following possible gap-closing scenario regarding retakes: i) female students may be more likely to take retake exams to improve their grade and/or ii) female students may make larger gains on retake exams. Examination of the data shows that the first possibility is wrong. We find that 78% $\pm$ 4% of male students retook at least one exam and 74% $\pm$ 3% of female students retook at least one exam. All together male students retook an average of 1.41 $\pm$ 0.09 exams and female students retook an average of 1.33 $\pm$ 0.07 exams. It seems that the two genders had about the same need for retake exams.
@@
We can examine the second possibility mentioned above by measuring the gender gap on each exam and also the gender gap after each retake had been scored. As noted in section III, the grades from the 12 exams from the three different courses were normalized so that i) each set of first-attempt scores from each exam in each course has the same average, = 0, and standard deviation, = 1, and ii) the recorded grades (i.e. after retake) from each exam from each class were similarly z-scored. The result of this is that each gender gap will be measured in units of standard deviations of the exam scores for that exam. We will also control for each student’s incoming GPA in our estimates of each in-term exam gender gap. We do this because the exam-score gender gap in a class has a small but clear dependence on the GPA gender gap in that class (see Appendix A which shows these class-dependent gender gap data) and we wish to remove that dependence from our in-term exam results. We control for GPA using hierarchical linear modeling (HLM) with the students as lowest level in the hierarchy, the set of classes as the next higher level, and then the part of the 3-course series (A, B, or C) makes up highest level of the hierarchy. We use all three levels because our retake classes cover all three levels. We use the variables $GPA$ for a student’s GPA and $Female$ = 1 for female students and = 0 for male students and fit the following model eight times, once for each of the four first-attempt exam scores taken through the term and once for each of the four exam scores recorded after the retake scores are factored in,
@@ IV.2 Low-stakes versus medium-stakes assessments
$b_{GPA}$ 0.8121 0.0047 173.0 $<10^{-3}$
$b_{Female}$0.2262 0.0046 -49.0 $<10^{-3}$
$b_{Retake}$0.169 0.052 -3.24 0.001
$b_{Female*Retake}$ 0.234 0.057 4.10 $<10^{-3}$
$b_{0}$2.310 0.018 -128.1 $<10^{-3}$
IV.3 Highest stakes assessments
Since 203 of the classes in our database have final exam scores it seems useful to examine those data because the final exam scores contribute much more to the course grade than any in-term exam and so can be considered to be high-stakes exams. Unfortunately we don’t have a low-stakes version of the final exam to compare with the high-stakes versions because even in the courses with retake options on the exams, it was not possible to retake the final exam. However, we do have the medium-stakes in-term exam scores in each class which should be good predictors of final exam scores in that class because they are written and graded by the same instructors and largely cover the same material. In fact, we find a moderate/strong correlation between the in-term exam average and the final exam score (correlation coefficient = 0.65, $p<10^{-3}$) but a smaller correlation between student GPA and either in-term or final exam score (both correlation coefficients are about 0.56, $p<10^{-3}$). We’ve already suggested that the medium-stakes in-term exam scores include a bias against female students, so it may be that using these scores as predictors of final exam scores will not only control for the demonstrable knowledge/ability of each student in the particular course but may also control for any gender bias inherent in exams. If this were the case then the controlled final exam gender gap should be approximately zero. To test this idea we will use HLM to fit the following model
@@ IV.3 Highest stakes assessments
The resulting coefficients from the HLM fit to Eq. 3 are given in Table 3. The most notable coefficient is $b_{Female}$ and the most notable thing about this coefficient is that it is quite clearly not zero. In fact, the (in-term exam controlled) final exam gender gap is half the size of the gender gap for the in-term exams alone. This tells us that the full gender gap on the high-stakes final exam is about 50% larger than that of the medium-stakes in-term exams. So the high-stakes final exam has a much larger gender gap than the medium-stakes in-term exams which, themselves, have a much larger gender gap than the low-stakes first-attempt in-term exams in the retake classes. And $b_{Female*Retake}$ is consistent with there being a similar (in-term-exam controlled) gender gap in the high-stakes final exam in the retake classes. Again, it seems likely that the importance (or weight) of the exam (the exam stakes) is closely related to the size of the grade penalty received by female students. Finally, as noted in Appendix B, we have checked that these coefficients and the resulting conclusions hold when we (randomly) include each student only once in our analyses.
@@ VI Limitations
All of the retake exam classes (as well as the non-retake classes) were offered within the CLASP (Collaborative Learning through Active Sense-making in Physics) series of courses [35, 33]. These courses are studio type courses where students spend almost 80% of their class time working in discussion/lab sections, generally in groups of five. The effects of allowing retake exams may well be different in classes that are predominantly lecture-based. In addition, almost the entire course grade in the courses studied was determined by the exam grades in these courses. It could be that in a course where most of the grade is determined by timed in-person solo assessments, differences of the impact of the stakes of each of those assessments are easier to identify. We encourage replication studies of this phenomenon.
@@ Appendix B Including a student only once doesn’t change our numbers much
$b_{GPA}$ 0.8052 0.0072 112.0 $<10^{-3}$
$b_{Female}$0.2226 0.0072 -31.0 $<10^{-3}$
$b_{Retake}$0.164 0.055 -3.00 0.003
$b_{Female*Retake}$ 0.233 0.061 3.85 $<10^{-3}$

### v3

# High stakes exams inflate a gender gap and contribute to systematic grading errors in one introductory physics series

@@ I.1.1 Gaps - student level characteristics
Many different explanations for the existence of these types of gender gaps have been proposed and investigated. Some of the leading theories for the gender gaps in physics are that women on average i) have poorer physics and/or math preparation [9, 10, 6], ii) have less overall interest in physics [11], and/or iii) suffer more from test anxiety, which is amplified on high-stakes exams [12, 13, 14, 10, 15, 16, 7]. It’s also been shown that women don’t feel like they belong in physics as much as their male peers [17, 4], that they have less of a science or physics identity [18], and/or have lower self-efficacy in physics [10, 15, 19, 20]. Finally, the context (e.g. physical vs biological or practical vs theoretical) of an exam question may influence [21, 22] the gender gap for that question.
@@ I.1.2 Gaps - system level issues
Leslie et al. [24] take a more systemic view in noting that gender gaps in representation are largest in a few specific academic fields (including physics) where the practitioners in the field are more likely to think that to succeed in their field “hard work alone just won’t cut it; you need to have an innate gift or talent.” Muradoglu et al. [25] use this idea, that success in a field is perceived to require innate brilliance, together with the general societal view that women are not as intellectually talented as men [26], to argue that a natural result is that, on average, women would have lower average self-efficacy in such fields. Since, self-efficacy inversely correlates with anxiety, it also is reasonable that women, on average, would be expected to be more anxious than men during the high stakes exams in these particular academic fields. As noted earlier, in this paper we will show data consistent with the idea that there is no gender gap in physics understanding/ability but that the commonly measured exam gender gap arises because of the stakes associated with that exam. In this sense the exam gender gap may simply be an incorrect measurement of the underlying knowledge/ability. It would be ironic if such illusory exam gender grade gaps were part of the evidence society used to conclude that women were not as talented as men at physics and that that conclusion led to women being more anxious on, and under-performing on, exactly these high stakes exams that led to society’s conclusions.
@@ I.2 Retake Exams in Physics
One possible way to lower the stakes of an exam, while still using exams as an assessment, is to offer a retake option for each exam given during the term. For each in-term exam, a retake covers the same topics as the first-attempt exam, is offered after the students have seen their score on their first-attempt, is offered to every student regardless of grade, and is given a score that can supplant the first-attempt score. Thus far, there is very little research on retake exams in large introductory STEM classes. Davis et al. [28] find that students in a large introductory human physiology course who took advantage of retake exams on average increase their grade by 6%. Kortemeyer et al. [29] share the benefits of retakes as learning tools in introductory physics. However, none of the examples we found examined the relationship between the retakes and the gender gap in physics courses aside from our own previous work on this same dataset.
@@ III Database Details
All of the exam data we discuss here are from classes in the introductory physics series taken mostly by bioscience students and offered between 1997 and 2015 at an R1 public university. These dates were chosen because these years included the retake classes, and because the classes that allowed students to take retake exams are part of the same course series as those that did not allow retakes. The topics covered in the course series did not change over the set of years included in the study and most of the specific student activities are also essentially the same over those years. We chose this set of years because we had access to individual exam scores for almost all of those years but for no later years. We have access to individual exam scores (both first-take and retake scores) for three of the four classes that offered students retake exams. In our previous paper [1] we agued that the retake effect appeared to be independent of instructor because these instructors also taught non-retake classes and their course grades showed clear gender gaps in those classes. Finally, exams in these courses are open-ended in the sense that the students are expected to provide both a written discussion of some particular problem as well as calculate some result(s). All of the exams given in the retake courses were drawn from a pool of exams that had been given previously in these courses so we consider them to be representative of usual exams in these classes.
To compare the exam scores from the retake classes to the exam scores from the regular classes which did not offer retakes we add, to the database, exam data of the 254 non-retake classes from the same series for which we have scores from exams given to students during the term. We also added each student’s gender and race/ethnicity as well as each student’s university GPA upon entering the class. These data were obtained from the university administration and included before anonymizing the entire dataset several years ago. Since the numerical exam scores are given in a wide variety of ways by the many instructors of these classes, we have z-scored each exam in each course to give each exam grade distribution the same average, 0, and the same standard deviation, 1.
@@ IV.1 In-term exams in retake classes
Our expectation was that students who had high scores on their first attempt would not risk lowering that score and so would not take a retake exam. We found that on average a first attempt of an exam was followed by 34% of the students taking the retake a week later. The retake classes were anomalous in having zero gender gap in course grades (i.e. female students had the same average course grades as male students unlike in the regular courses) suggesting the following possible gap-closing scenario regarding retakes: i) female students may be more likely to take retake exams to improve their grade and/or ii) female students may make larger gains on retake exams. Examination of the data shows that the first possibility is wrong. We find that 78% $\pm$ 4% of male students retook at least one exam and 74% $\pm$ 3% of female students retook at least one exam (a $\chi^{2}$ test showed $\chi^{2}(1,N=399)=0.89$ with $p=0.35$ so no measurable gender difference). All together male students retook an average of 1.41 $\pm$ 0.09 exams and female students retook an average of 1.33 $\pm$ 0.07 exams (a t-test gives $p=0.49$ so no difference here either). It seems that the two genders had about the same need for retake exams.
@@
We can examine the second possibility mentioned above by measuring the gender gap on each exam and also the gender gap after each retake had been scored. As noted in section III, the grades from the 12 exams from the three different courses were normalized so that i) each set of first-attempt scores from each exam in each course has the same average, = 0, and standard deviation, = 1, and ii) the recorded grades (i.e. after retake) from each exam from each class were similarly z-scored. The result of this is that each gender gap will be measured in units of standard deviations of the exam scores for that exam. We will also control for each student’s incoming GPA in our estimates of each in-term exam gender gap. We do this because the exam-score gender gap in a class has a small but clear dependence on the GPA gender gap in that class (see Appendix A which shows these class-dependent gender gap data) and we wish to remove that dependence from our in-term exam results. We control for GPA using hierarchical linear modeling (HLM) with the students as lowest level in the hierarchy, the set of classes as the next higher level, and then the part of the 3-course series (A, B, or C) makes up highest level of the hierarchy. We use all three levels because our retake classes cover all three levels. With HLM the data are fit first class by class and then those fits are combined into A, B, and C results which are then combined to give the final coefficients and their errors. We use HLM to account for class-to-class differences in exams and A, B, and C differences in topics covered since any of these may be affecting the gender gaps and we want the final error estimates to reflect all of these variations. For more detailed information on HLM see reference [34]. We use the variables $GPA$ for a student’s GPA and $Female$ = 1 for female students and = 0 for male students and fit the following model eight times, once for each of the four first-attempt exam scores taken through the term and once for each of the four exam scores recorded after the retake scores are factored in,
@@ IV.2 Low-stakes versus medium-stakes assessments
$b_{GPA}$ 0.8121 0.0047 173.0 $10^{-3}$
$b_{Female}$ -0.2262 0.0046 -49.0 $10^{-3}$
$b_{Retake}$ -0.169 0.052 -3.24 0.001
$b_{Female*Retake}$ 0.234 0.057 4.10 $10^{-3}$
$b_{0}$ -2.310 0.018 -128.1 $10^{-3}$
IV.3 Testing the stakes hypothesis using the highest stakes assessments
Since 203 of the classes in our database have final exam scores it seems useful to examine those data because the final exam scores contribute much more to the course grade than any in-term exam and so can be considered to be high-stakes exams. Unfortunately we don’t have a low-stakes version of the final exam to compare with the high-stakes versions because even in the courses with retake options on the exams, it was not possible to retake the final exam. However, we do have the medium-stakes in-term exam scores in each class which should be good predictors of final exam scores in that class because they are written and graded by the same instructors and largely cover the same material. In fact, we find a correlation coefficient between the in-term exam average and the final exam score of 0.65 (with $p<10^{-3}$) but a smaller correlation between student GPA and either in-term or final exam score (both correlation coefficients are about 0.56, $p<10^{-3}$). We’ve already suggested that the medium-stakes in-term exam scores include a bias against female students, so it may be that using these scores as predictors of final exam scores will not only control for the demonstrable knowledge/ability of each student in the particular course but may also control for any gender bias inherent in exams. If this were the case then the controlled final exam gender gap should be approximately zero. To test this idea we will use HLM to fit the following model
@@ IV.3 Testing the stakes hypothesis using the highest stakes assessments
The resulting coefficients from the HLM fit to Eq. 3 are given in Table 3. The most notable coefficient is $b_{Female}$ and the most notable thing about this coefficient is that it is quite clearly not zero. In fact, the (in-term exam controlled) final exam gender gap is half the size of the gender gap for the in-term exams alone. This tells us that the full gender gap on the high-stakes final exam is about 50% larger than that of the medium-stakes in-term exams. So the high-stakes final exam has a much larger gender gap than the medium-stakes in-term exams which, themselves, have a much larger gender gap than the low-stakes first-attempt in-term exams in the retake classes. And $b_{Female*Retake}$ is consistent with there being a similar (in-term-exam controlled) gender gap in the high-stakes final exam in the retake classes. Again, it seems likely that the importance (or weight) of the exam (the exam stakes) is closely related to the size of the grade penalty received by female students. Finally, as noted in Appendix B, we have checked that these coefficients and the resulting conclusions hold when we (randomly) include each student only once in our analyses. Again, as in section IV.2, it seems that exams with higher stakes have higher gender gaps.
@@ VI Limitations
All of the retake exam classes (as well as the non-retake classes) were offered within the CLASP (Collaborative Learning through Active Sense-making in Physics) series of courses [38, 36]. These courses are studio type courses where students spend almost 80% of their class time working in discussion/lab sections, generally in groups of five. The effects of allowing retake exams may well be different in classes that are predominantly lecture-based. In addition, almost the entire course grade in the courses studied was determined by the exam grades in these courses. It could be that in a course where most of the grade is determined by timed in-person solo assessments, differences of the impact of the stakes of each of those assessments are easier to identify. The data collected in this study are at least a decade old. Since that time, there have been significant changes to the educational landscape including advances in AI, a worldwide pandemic, and a changing student population. These and other factors may affect future results. Future research could compare exams from courses with large and small (or reverse) gender gaps to further understand how points are allocated on exams with different stakes. We encourage replication studies of the exam retake and stakes phenomena.
@@ Appendix B Including a student only once doesn’t change our numbers much
$b_{GPA}$ 0.8052 0.0072 112.0 $10^{-3}$
$b_{Female}$ -0.2226 0.0072 -31.0 $10^{-3}$
$b_{Retake}$ -0.164 0.055 -3.00 0.003
$b_{Female*Retake}$ 0.233 0.061 3.85 $10^{-3}$